OpenStudy (anonymous):

(x-2)^2=36, give the exact and approximate solutions to three decimal places. x=?

OpenStudy (anonymous):

x^2+4=36 x^2=36-4 x^2=32 x=4\[\sqrt{2}\]

OpenStudy (anonymous):

x=4squareroot of2